OpenStudy (anonymous):

The graph of y = ax^2 + bx + c is shown below. Determine the solution set of 0 = ax^2 + bx + c.

OpenStudy (anonymous):

\[\left\{x | x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a} | \sqrt{b^2 - 4ac} \ge 0 \right\}\]